OpenStudy (e.cociuba):

What is the molarity of a solution containing 23g of NaCl in 500 mL of Solution? (The molecular mass of NaCl is 58.44.) I need some help solving this. Thanks! :) a.) 46 M b.) 2.7 M c.) 0.79 M d.) 0.50 M

OpenStudy (mertsj):

What is the molecular weight of NaCl?

OpenStudy (mertsj):

Oh. Sorry. I see it now. It is posted.

OpenStudy (e.cociuba):

Ya :)

OpenStudy (e.cociuba):

Do I use this formula: ??

OpenStudy (mertsj):

\[\frac{23 g NaCl}{500 ml solution}\times \frac{1 mole NaCl}{58.44 g NaCl}\times\frac{1000 ml}{1L}\]

OpenStudy (mertsj):

Yes. You could use the formula you posted but you would have to change the grams to moles first because that formula says that molarity = number of moles/volume

OpenStudy (e.cociuba):

 I need help on wat to do after tht...